10) Consider an industry with a linear inverse demand, p = 100 - Q, and MC = AC = $10. Solve for industry output, price, and profits if the industry is:

  1. Perfectly competitive

  1. Monopolistic
Answer #1

10) Under perfect competition; equilibrium condition is P = MC

100 - Q = 10

Q = 90 units

P = 100 - 90 = $ 10

Profit = TR - TC = PQ - Q x AC = 10 x 90 - 90 x 10 = 900 - 900

Profit = 0

Under monopolistic; Equilibrium condition is MR = MC

TR = PQ = (100 - Q)Q = 100Q - Q2

MR = 100 - 2Q

MC = 10

100 - 2Q = 10

2Q = 90

Q = 45 units

P = 100 - 45 = $ 55

Profit = TR - TC = 55 x 45 - 55 x 10 = 2475 - 550

Profit = $ 1925